You are here: Foswiki>AdaptiveOptics Web>ADSECDXDetach (08 Jun 2013, JuanGuerra)Edit Attach


AdsecDX detached.

Arcetri inspection.


Staff: Armando riccardi (Arcetri),Juan Carlos Guerra (LBTO),Michael Wagner (LBTO)

First inspection of the thin shell after it was removed last Friday,May31.

Armanod report;

The fraction of coating that apparently went off around those magnets is about 50%. We tested electrical continuity between outside the apparently stripped area and inside it with a Ohm-meter giving an open circuit reading. That is a DC measurement and tomorrow morning we will test also in AC injecting the reference square wave signal on the electrical coating of the shell and see if the AC impedance will reduce.


Starting meeting @ 8:30 am; Armando riccardi (Arcetri),Juan Carlos Guerra (LBTO),Michael Wagner (LBTO) and Michael Lefvre(LBTO)

Setting up equipment for the electronic test of the back of the shell using the reference signal.

- Measurement of the reference signal. - Inspection of the distribution boards.

Summary reported by Armando.

TS3 is heavily affected by the loss of coating in the area of capsens. In particular in the outer ring of actuators, there is a bad clustering of actuators with large fraction of stripped coating. I suggest the use of TS4 to avoid a severe penalty of the performances of the mirror unit. We still working on cleaning TS3 that can be used to further test the electronics while waiting for TS4. We have a spare problem with the distribution board that have been affected by contamination, we will try to clean and fix when spare will not available.

There is a link to a document with more details in: Also the document is in the attachment.


Starting meeting @ 8:30 am; Armando Riccardi (Arcetri),Juan Carlos Guerra (LBTO) and Michael Wagner (LBTO)

- remove the distribution boards, those are damage or with signal of contamination. - clean and using ultrasonic bath. - start to clean the back of the shell; acetone. - remove the bad actuators with a signal of rust on the front.

The distribution boards started to be cleaning with the ultrasonic bath and it looks like that they can be clean without any harsh.

An inspection of the holes in the reference body shows that they are many holes that are contaminated, rusted or completely gone.

Armando has sent the report and this can be found in : The document is also upload in this wiki.


Starting meeting @ 8:30 am; Armando Riccardi (Arcetri),Juan Carlos Guerra (LBTO) and Michael Wagner (LBTO) Call to Roberto Biassi (Microgate) to discuss the issue found in the actuators holes, where it has been found many rust and degraded of the contacts. This requires an extra work and it is needed to dismount all the actuators.

With the number of actuators that we have removed we can not guarantee 100% of that the others actuators holes are ok.

It has been arrange another call to Roberto after Julian arrives to the mountain.

Dismounting and mapping the bad holes.

At 18:00 all the actuators were dismount and all the distributions boards were cleaning. Actuators ready for cleaning. Also the bad holes will be cleaning


Preparing to clean the holes and the actuators.

-- JuanGuerra - 05 Jun 2013

  • measurement of the square signal:

I Attachment Action Size Date Who Comment
P6046672.JPGJPG P6046672.JPG manage 1 MB 06 Jun 2013 - 16:00 UnknownUser measurement of the square signal
unit_rescue_v2.docxdocx unit_rescue_v2.docx manage 6 MB 06 Jun 2013 - 15:57 UnknownUser document
unit_rescue_v3.docxdocx unit_rescue_v3.docx manage 9 MB 07 Jun 2013 - 15:35 UnknownUser summary
Topic revision: r6 - 08 Jun 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback